OpenStudy (anonymous):

You borrow $500 from your parents to buy a new computer. Your parents charge you 3% simple interest and give you two years to repay the money How much interest do you pay on the loan??

OpenStudy (anonymous):

\[S.I = \frac{ P \times R \times T }{ 100 }\] S.I=Simple Interest P=Principal R=Rate T=Time